Calculate the time it will take to deposit 1 mole of silver if a current of 6A is passed through a solution of silver trioxonitrate (v) (F=96500)
1
Expert's answer
2020-11-04T14:18:39-0500

Silver trioxonitrate, or silver nitrate AgNO3 solution electrolysis reaction at cathode is the following:

Ag+ + e- "\rightarrow" Ag.

According to the Faraday's laws of electrolysis, the number of the moles of the deposited silver is proportional to the charge passed through the solution:

"n(Ag) = \frac{It}{F}" .

Therefore, the time needed to deposit 1 moles of silver with a current of 6 A is:

"t = \frac{n(Ag)·F}{I} = \frac{1 \text{ mol}·96500\text{ C/mol}}{6\text{ A}} = 16083" seconds, or 268 minutes, or 4 hours 28 minutes.

Answer: it will take 4 hours 28 minutes to deposit 1 mole of silver if a current of 6 A is passed through a solution of silver trioxonitrate.